Grade 11Mechanics

A stone is projected with a velocity 20√2 m/s at an angle of 45° to the horizontal .The average velocity of stone during the motion from starting point to its maximum height is

sanjana warma
5 Years agoGrade 11

1 Answer

Vikas TU
5 Years ago
Dear student 
Displacement at highest point will be (R/2,H) in horizontal and vertical direction respectively.
Average velocity = Displacement /Time 
V = 20sqrt(2) 
Angle Q = 45 degree 
H= V^2Sin^2Q/2g 
R = V^2Sin^2Q/g
Time = VsinQ /g 
H = 20 m 
R/2 = 40m 
T = 2 
Average velocity = 10sqrt5 m/s